Find two consecutive positve integers such that the sum of thier square is 85.
Answer by checkley71(8403)

x^2+(x+1)^2=85
x^2+x^2+2x+1-85=0
2x^2+2x-84=0
(2+14)(x-6)=0
x-6=0
x=6 answer for the smaller number
6+1=7 the larger number.
proof
6^2+7^2=85
36+49=85
85=85